OpenStudy (anonymous):

ALgebra 2 help?! (3x+12)/6

OpenStudy (eric_d):



OpenStudy (igreen):

(3x + 12)/6 Divide: 1/2x + 2

OpenStudy (igreen):

3 / 6 = 1/2 12 / 6 = 2

OpenStudy (anonymous):